OpenStudy (anonymous):

In a video game, minerals are mined at a rate of 0.8 minerals per second. The player starts off with 250 minerals. This can be represented by M(s) = 250 +0.8s. What does M(45) represent? The time it takes to mine 45 minerals The number of minerals mined in 45 seconds The time it takes to mine 250 minerals None of these I cant decided if it's b or d? help please.

OpenStudy (doc.brown):

A function takes the crap from inside the brackets and sticks it into the formula. In this case the s. So \[M(s)=250+0.8s\]\[M(45)=250+0.8(45)\]\[M(45)=250+36\]\[M(45)=268\]Which is the starting inventory plus the number of minerals mined in 45 seconds. Answer d.

OpenStudy (anonymous):

Thank you (:

OpenStudy (doc.brown):

The time it takes to mine M(45) or M(250) minerals:\[M(m)=\frac{m}{0.8}\]The number of minerals mined in M(45) seconds\[M(s)=0.8s\]